mathslover (mathslover):

At what height will a man's weight be half his weight as on the surface of earth? (here R is the radius of earth) A) 0.236 R B) 0.732 R C) 0.414 R D) 0.645 R

mathslover (mathslover):

@mukushla

OpenStudy (anonymous):

\[\frac{m_2g_2}{m_1g_1}=\frac{1}{2}\]

mathslover (mathslover):

hmn wait...

mathslover (mathslover):

so I got this...

mathslover (mathslover):

m i right or m i right? :P well any mistake there " mukushla "sir""

OpenStudy (anonymous):

quite right

mathslover (mathslover):

Oh k thanks for that sir... @mukushla

OpenStudy (anonymous):

no problem kush

OpenStudy (anonymous):

acceleration due to gravity at height h from the surface of the earth= (GM/R^2) (1+h/r) ^(-2)

mathslover (mathslover):

I got it @akash123 thanks

OpenStudy (anonymous):

what's the answer....if u dont neglect the higher order terms...if i neglect higher order terms then h=0.25R...but it's not in the options

mathslover (mathslover):

h = 0.414 R

OpenStudy (anonymous):

so u calculated the sum of series...s=1-2x+3x^2-4x^3...

OpenStudy (anonymous):

got it..i was on d wrong track

mathslover (mathslover):

no problem